John is building a rectangular chicken coop in his backyard. He plans to use the existing fence as one side of the coop. John has 37 Feet of fencing.
Create a function that represents the area of the chicken coop.
What are the dimensions of the chicken coop that maximizes the area? What is the maximum area of the chicken coop?

Answer by stanbon(75887)

John is building a rectangular chicken coop in his backyard. He plans to use the existing fence as one side of the coop. John has 37 Feet of fencing.
Create a function that represents the area of the chicken coop.
What are the dimensions of the chicken coop that maximizes the area? What is the maximum area of the chicken coop?
-----
Note:: Need the fencing for 3 sides of the rectangle
Let width = w
Then length = 37-2w
--------------
Area = width*length = -2w^2 + 37w
--------------------
Max area occurs when w = -b/(2a) = -37/(2*-2) = 9.25
Then length = 37-2w = 37-18.5 = 18.5 ft
------
Max Area = W*L = 9.25*18.5 = 171.125 sq. ft
-------------
Cheers,
Stan H.
-----------